# People who frequently use ChatGPT for writing tasks are accurate and robust detectors of AI-generated text

### v1

@@ D.1 Categorization of Comments
Table 17: Truncated prompt used for comment analysis. The first insert is filled by a list of definitions of the categories found in Table 4 and the second insert is filled by an annotator explanation. A full version of the guide can be found at https://github.com/jenna-russell/human_detectors.
D.2 Individual Experts Commentary
Each expert had clues they favored using throughout all experiments. Annotator 1,whose category mention frequencies can be found in Figure 9, Figure 10 depicts Annotator 2 comments, Figure 11 shows Annotator 3 comments, Figure 12 shows Annotator 4 comments and Figure 13 has commentary frequencies from Annotator 5. The individual heatmaps highlight the range of clues used by annotators, who individually had no reference of clues other experts used.
@@ Appendix E Automatic Detection
E.1 Explainable Detection
In this section we detail the prompt based detector used for experiments in §4.

### v2

@@ D.1 Categorization of Comments
Category Freq Definition Example Explanations
Table 17: Full taxonomy of clues used by experts to explain their detection decisions. For each category, we report the frequency of explanations that mention that category (regardless of if the annotator was correct) and provide examples of explanations for both human-written and AI-generated articles. While vocabulary and sentence structure form the most frequent clues, more complex phenomena like originality, clarity, formality, and factuality are also distinguishing features.
@@
Table 18: Truncated prompt used for comment analysis. The first insert is filled by a list of definitions of the categories found in Table 3 and the second insert is filled by an annotator explanation.
D.2 What do expert annotators see that nonexperts don’t?
Figure 9: Same as Figure 3, except only computed over explanations when experts were incorrect. Formality is a big source of misdirection for o1-Pro articles, while fixating on sentence structure can lead experts to false positives. Details of each category can be found in Table 3.
D.3 Individual Experts Commentary
Each expert had clues they favored using throughout all experiments. Annotator 1,whose category mention frequencies can be found in Figure 10, Figure 11 depicts Annotator 2 comments, Figure 12 shows Annotator 3 comments, Figure 13 shows Annotator 4 comments and Figure 14 has commentary frequencies from Annotator 5. The individual heatmaps highlight the range of clues used by annotators, who individually had no reference of clues other experts used. We again note that our experts are untrained at this detection task, and they could likely improve their individual performance if provided with feedback on their errors.
@@ D.3 Individual Experts Commentary
Annotators don’t always focus on the same clues:
Our experts focus on different properties of the text to arrive at their decisions. Annotator 1 is the only one to pick up on “AI names”: in fact, 63.3% of GPT-4o and 70% of Claude-3.5-Sonnet articles include either the name Emily or Sarah.313131o1-Pro favors names of real people instead of fictional names when generating articles. Annotator 2 and 3 emphasize linguistic features, such as humans not always following “proper” writing conventions or the tendency of LLMs to always list examples in groupings of three. Annotator 4 focuses on the flow of the articles, analyzing the specificity of detail and motivation behind an article, while Annotator 5 frequently mentions whether a quotation sounds natural.
Commentary on Paraphrasing Attack in Experiment 3:
Annotators continue to pick up on many of the same clues within the paraphrased articles that were also apparent in Experiments 1 & 2, such as high frequency of “AI vocab” (even after paraphrasing), formulaic sentence structures, and cheerful summary conclusions. Somewhat counterintuitively, explanations about paraphrased articles note AI vocab in 88% of explanations, compared to only 69.8% of non-paraphrased GPT-4o articles. Similarly, quotations are mentioned in 33.8% of explanations about paraphrased articles, a much higher rate than other configurations: a close reading of explanations reveals that experts flagged quotes that were always in the same format and style (e.g., only placed at the end of each paragraph).
One Annotator Finds o1 Content Hard to Detect
Interestingly, Annotator 3’s TPR drops considerably, as they prioritize signs of human writing over signs of AI writing when making their judgments. Annotator 3, the only one consistently fooled by o1-Pro outputs, relied too much on signs of informality (e.g., contractions, slang usage, usage of just and actually) as a sign of human writing, with 66.7% of their explanations relating to formality (Figure 12).
@@ Appendix E Automatic Detection
E.1 Automatic Detectors Benchmarked
In this section, we detail the detectors used as automatic benchmarks in Table 2.
Thresholding for Automatic Detectors:
E.2 Explainable Detection
In this section we detail the prompt based detector used for experiments in §4. While humans experts are found to be robust detectors, an obvious drawback is that hiring humans is expensive and slow: on average, we paid \$2.82 per article including bonuses, and gave annotators a week to complete each experiment. We wondered if we could prompt an LLM to do the same.